# TeaMs-RL: Teaching LLMs to Generate Better Instruction Datasets via Reinforcement Learning

### v3

@@ 4.4 Experiments of Model Privacy Attack
4.5 Comparison Experiments on Apacal Eval Benchmarks
@@ 4.6 Comparison Experiments of Solving General Tasks
In this evaluation, we present scenarios where random tasks are provided, and the subsequent response is generated by our model (TeaMs-RL-1-7b) and baseline models. Notably, our model performs exceedingly better than the baseline models in providing a well-structured and accurate response to the given instruction. More cases are provided in Appendix F and Appendix G.
@@ 4.7 Comparison Experiments of Solving a Math Problem
We conduct comparison experiments with several strong baseline models on mathematical tasks. The experimental procedure entails inputting instructions into LLMs, with the LLMs subsequently generating responses aligned with the provided instructions. The empirical findings from our study reveal that the post-alignment LLM, TeaMs-RL-7b, exhibits superior mathematical capabilities compared to the baseline models. This enhanced performance is noteworthy, particularly when it surpasses even those models with a larger parameter count, such as the 13b models. This outcome underscores the efficacy of our alignment and training methodologies in augmenting the specific skill set of LLMs, particularly in mathematical problem-solving. More results are introduced in Appendix H.
@@ 4.9 Ablation Experiments Regarding Data size
To further assess the efficacy of our proposed method, we conduct ablation experiments focusing on the impact of data size, as shown in Table 3. By expanding the dataset from 17,878 to 19,395 entries, notable improvements are observed in the performance of the TeaMs-RL-7b model on both the Hellaswag and ARC Challenge benchmarks. Specifically, the model’s performance on the Hellaswag benchmark increases from 77.17 to 78.35, while on the ARC Challenge, it increases from 54.35 to 55.89. These results provide further confirmation of our method’s effectiveness.
@@
Table 3: Ablation experiments regarding the impact of data size: expanding the dataset from 17,878 to 19,395 entries, the performance of the TeaMs-RL-7b model improved on both the Hellaswag and ARC Challenge benchmarks, with scores rising from 77.17 to 78.35 and 54.35 to 55.89 respectively.
@@ 6 Limitations
Appendix
[…]
1
Content
Appendix A Practical Algorithm
Algorithm 2 Teams-RL: Teaching LLMs to Teach Themselves Better Instructions via RL.
@@ Appendix A Practical Algorithm
6:     Use the methods of conjugate gradient and line search to search the policy parameters $\theta_{k+1}$.
@@
Appendix B Action Set
@@
Appendix C Generate Imaginative Instructions by Teaching ChatGPT-4
C.1 Instruction 1
@@ C.8 Instruction 8
Appendix D Generate Imaginative Instructions by Teaching ChatGPT-3.5
D.1 Instruction 1
@@ D.6 Instruction 6
Appendix E More Comparison Results between WizardLM and TeaMs-RL
@@
Appendix F More Comparison Experiments of Solving Chat Tasks
F.1 Task 1
@@ F.4 Task 4
Appendix G TeaMs-RL-7b Performance on General Tasks
G.1 Task 1
@@ G.3 Task 3
Appendix H More Comparison Experiments of Solving a Math Problem
@@ H.4 Task 4
Appendix I Experiment Settings

### v4

@@ 4.3 Model Privacy Attack Experiments
4.4 Comparison Experiments on Alpaca Eval Benchmarks
@@ 4.5 Comparison Experiments of Solving General Tasks
In this evaluation, we present scenarios where random tasks are provided, and the subsequent response is generated by our model (TeaMs-RL-1-7b) and baseline models. Notably, our model performs exceedingly better than the baseline models in providing a well-structured and accurate response to the given instruction. More cases are provided in Appendix H and Appendix I.
@@ 4.6 Comparison Experiments of Solving a Math Problem
We conduct comparison experiments with several strong baseline models on mathematical tasks. The experimental procedure entails inputting instructions into LLMs, with the LLMs subsequently generating responses aligned with the provided instructions. The empirical findings from our study reveal that the post-alignment LLM, TeaMs-RL-7b, exhibits superior mathematical capabilities compared to the baseline models. This enhanced performance is noteworthy, particularly when it surpasses even those models with a larger parameter count, such as the 13b models. This outcome underscores the efficacy of our alignment and training methodologies in augmenting the specific skill set of LLMs, particularly in mathematical problem-solving. More results are introduced in Appendix J.
@@ 4.8 Ablation Experiments Regarding Data size
To further evaluate the effectiveness of our proposed method, we conduct ablation experiments to analyze the impact of dataset size, as summarized in Table 3. Expanding the dataset from 17,878 to 19,395 entries led to notable performance improvements for the TeaMs-RL-7b model on both the Hellaswag and ARC Challenge benchmarks. Specifically, the model’s performance on the Hellaswag benchmark improves from 77.17 to 78.35, while on the ARC Challenge benchmark, it increases from 54.35 to 55.89. Further expanding the dataset from 19,395 to 21,396 entries yields more improvements: the Hellaswag benchmark score increases from 78.35 to 78.59, and the ARC Challenge score improves from 55.89 to 56.74. These results provide further confirmation of our method’s effectiveness. It is worth noting that the performance improvements are not strictly linear with the increase in dataset size; however, adding more data generally proves beneficial for enhancing model performance.
@@
21,396 56.74 78.59
Table 3: Ablation experiments regarding the impact of data size: Expanding the dataset from 17,878 to 21,396 entries improves the performance of the TeaMs-RL-7b model on both the Hellaswag and ARC Challenge benchmarks.
@@ 6 Limitations
Appendix A A Methodological Example
To clarify our methodology and facilitate replication, we provide a detailed example illustrating the inputs and outputs at each stage of our process, similar to Figure 1 in our paper.
Stage 1: Policy Training
Components:
  • RL Policy: (derived from Instructor Model) Learns to select actions based on the reviewer’s feedback to improve the instructions.
  • Reviewer Model: Provides feedback on the instructions.
Example:
  1. 1.
    RL Policy Training
    With any Initial Instructions such as:
    “Describe the process of photosynthesis.”
    The RL policy with Instructor Models learns to select actions that diversify instructions based on feedback from the reviewer model.
  2. 2.
    RL Policy Training:
    Based on the reviewer’s feedback, the RL policy learns to select actions such as:
    • Add Constraints
    • Deep Reasoning
    • Width Reasoning
Stage 2: RL Policy Action Selection
Selected Action: "Add Constraints"
Purpose: To make the instruction more challenging and comprehensive by adding specific constraints or requirements.
Stage 3: Guiding Expert LLMs
Purpose: To make the instruction more challenging and comprehensive by adding specific constraints or requirements.
RL Policy Action Selection such as: "Add Constraints"
The RL policy generates a specialized prompt based on this action to guide the expert LLM in rewriting the instruction.
Add Constraints: Action Prompt to Expert LLM:
“““ I want you to act as a Prompt Rewriter. Your objective is to rewrite the given prompt into a more complex version to make it more challenging for AI systems like ChatGPT and GPT-4. Ensure that the rewritten prompt remains reasonable, understandable, and suitable for human response. Do not omit any non-text parts such as tables or code in the given prompt. Do not repeat conditions or requirements in your response, and do not disclose your role. Provide only the rewritten prompt without any introduction or explanation. The new prompt should not exceed 2048 words. You should complicate the given prompt by adding one more constraint or requirement. Try not to make the rewritten prompt verbose; you can only add or replace 10 to 20 words in the given prompt. Do not include phrases like ’Given Prompt’ or ’Rewritten Prompt’ in your response. Given Prompt: "Describe the process of photosynthesis."
Stage 4: Expert LLM Generates Rewritten Instruction
Expert LLM Output:
“Describe the process of photosynthesis in plants and explain how it varies in different environmental conditions.”
Explanation:
  • Original Instruction: Simple and straightforward.
  • Rewritten Instruction: Adds the constraint of explaining variations in different environmental conditions, increasing complexity and depth.
  • Compliance with Guidelines: The expert LLM added approximately 12 words, adhering to the limit of adding or replacing 10 to 20 words.
Once we have the generation instructions, we use them to query the expert model for corresponding responses. After obtaining the responses, we have the final dataset (instructions and corresponding responses), which is then used to fine-tune foundation models.
Summary
This example demonstrates how our RL policy guides expert LLMs to generate more complex and high-quality instructions by selecting appropriate actions. The process ensures that the generated instructions are challenging yet reasonable, facilitating the creation of a valuable dataset for training advanced AI models.
Appendix B Practical Algorithm
Algorithm 2 TeaMs-RL: Teaching LLMs to Generate Better Instruction Datasets via RL.
@@ Appendix B Practical Algorithm
6:     Use the methods of conjugate gradient and line search to search the policy parameters $\theta_{k+1}$ via TRPO.
@@
Here, we introduce Trust Region Policy Optimization (TRPO) (Schulman et al., 2015), which is outlined in Algorithm 3. In this context, $\rho_{\pi}$ represents the discounted state visitation frequencies, while $\epsilon=\max_{s,a}\left|A_{\pi}(s,a)\right|$ denotes the maximum absolute advantage function value across all states and actions. The performance objective, $\eta(\pi)$, is defined as the expected cumulative reward over time, $\eta(\pi)=\mathbb{E}_{s_{0},a_{0},\dots}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t})\right]$, where $\gamma$ is the discount factor.
Algorithm 3 TRPO(Schulman et al., 2015)
Initialize $\pi_{0}$.
for $i=0,1,2,\dots$ until convergence do
     Compute all advantage values $A_{\pi_{i}}$.
     Solve the constrained optimization problem
$\pi_{i+1}=\arg\max_{\pi}\left[L_{\pi_{i}}(\pi)-CD_{\text{KL}}^{\text{max}}(\pi_{i},\pi)\right]$
     where $C=\frac{4\epsilon\gamma}{(1-\gamma)^{2}}$ and $L_{\pi_{i}}(\pi)=\eta(\pi_{i})+\sum_{s}\rho_{\pi_{i}}(s)\sum_{a}\pi(a|s)A_{\pi_{i}}$
end for
Appendix C Action Set
@@
Appendix D Teaching LLMs to Generate Instructions via Trained Policy
To teach expert LLMs in the generation of high-quality instructions, we employ an RL policy that we have trained. This policy prompts the expert LLMs, enabling them to produce high-quality instructions incrementally. This process is achieved by synergizing the initial instructions and ensuring the output enhances each step. This approach facilitates the expert LLMs in refining their instructional outputs, harnessing the RL policy for optimal effectiveness in instruction generation. The related cases are in subsections D.1 and D.2. It is noteworthy that these initial instructions are culled from a well-established and widely-recognized dataset, namely the Alpaca dataset [^8] https://github.com/tatsu-lab/alpaca_eval.git. The experiment settings are available in Appendix K.
D.1 Teaching ChatGPT-4 to Generate Imaginative Instructions
For more cases, please see Appendix E.
D.2 Teaching ChatGPT-3.5 to Generate Imaginative Instructions
An initial instruction: “Find a unique way to use a rubber band.”
For more cases, please see Appendix F.
Appendix E Generate Imaginative Instructions by Teaching ChatGPT-4
E.1 Instruction 1
@@ E.8 Instruction 8
Appendix F Generate Imaginative Instructions by Teaching ChatGPT-3.5
F.1 Instruction 1
@@ F.6 Instruction 6
Appendix G More Comparison Results between WizardLM and TeaMs-RL
@@
Appendix H More Comparison Experiments of Solving Chat Tasks
H.1 Task 1
@@ H.4 Task 4
Appendix I TeaMs-RL-7b Performance on General Tasks
I.1 Task 1
@@ I.3 Task 3
Appendix J More Comparison Experiments of Solving a Math Problem
@@
Appendix K Experiment Settings